Circular No. 8980

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2008 (216) 1 AND S/2008 (216) 2 F. Marchis, Carl Sagan Center at the SETI Institute and University of

California at Berkeley; and his collaborators, P. Descamps and J. Berthier, Institut de Mécanique Céleste et de Calcul des Ephémerides, Observatoire de Paris; and J. P. Emery, University of Tennessee at Knoxville, report on observations of (216) Kleopatra, a large M-type main-belt asteroid, recorded with the 10-m Keck II telescope and its adaptive-optics (AO) system between Sept. 19.27 and 19.50 UT, when the target was at r = 1.237 AU with a phase angle of 7^{o}.5. The direct images recorded with the NIRC2 camera reach an angular resolution up to 0′′.032 (corresponding to a spatial resolution of 29 km) with the J_{cont} filter (central wavelength at 1.21 μm), permitting confirmation that the primary has the “dog-bone” shape revealed by Ostro et al. (2000, Science 288, 836). The new AO data show additionally the presence of a ∼ 5-km-diameter companion at 0′′.72 (projected distance of 650 km) in p.a. 322^{o}, detected in every image recorded over the 5.5-hr baseline. Careful analysis of the last set of data indicates the presence of a second satellite, smaller (∼ 3 km) and located closer to the primary at an apparent distance of 380 km (0′′.42) in p.a. 333^{o}. With (87) Sylvia (cf. IAUC 8582) and (45) Eugenia (cf. IAUC 8817), this is the third multiple system with two small-km-sized satellites to be discovered in the main asteroid belt.

COMET P/2008 R1 (GARRADD) Improved, elliptical orbital elements for this comet (cf. IAUC 8969)

from MPEC 2008-S46:

T = 2008 July 25.200 TT  ω = 256.^{o}436
e = 0.34203  Ω = 52.018 }2000.0
q = 1.79261 AU  i = 15.899
a = 2.72444 AU n^{o} = 0.219174 P = 4.50 years

COMET 205P/GIACOBINI MPEC 2008-S47 contains astrometry and ephemerides for components

A, B, and C (cf. IAUC 8978), and also revised orbital elements (with nongravitational parameters) satisfying the primary component during 1896– 2008. Visual total-magnitude estimates for component A: Sept. 11.07 UT, 12.7 (J. J. Gonzalez, Palencia, Spain, 20-cm reflector); 16.96, 12.9 (M. Goiato, Araçatuba, Brazil, 22-cm reflector); 19.83, 12.2 (Gonzalez, León).

2008 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green